OpenStudy (anonymous):

x to the 9th power (x + 8) + 4 (x + 8) = I need the factor

myininaya (myininaya):

each term has the factor (x+8) in common

myininaya (myininaya):

(x+8)(1+4)=I

myininaya (myininaya):

(x+8)(5)=I 5(x+8)=I I=5(x+8)

myininaya (myininaya):

is that suppose to be x^9(x+8)+4(x+8)?

myininaya (myininaya):

\[(x+8)(x^9+4)\]